# Theory of periodic convolutional neural network

### v1

@@ 1 Introduction
In this work, we focus on a new architectural variant we call the periodic CNN. The key modification is the enforcement of periodic boundary conditions in convolutional operations. Such a structure arises naturally in many scientific and engineering applications, for example, when signals live on tori, in materials with repeating microstructures, or in spectral data analysis.
@@
This dichotomy yields both theoretical insights and practical implications. On the theory side, it provides one of the first precise separation results for CNN variants, showing exactly where periodicity enables expressivity and where it constrains it. On the practical side, it suggests that periodic CNNs are particularly well-matched to high-dimensional structured signals where information is concentrated along d1d-1 directions, a common scenario in physics-informed learning and scientific computing.
@@
Ridge functions $f(x)=\phi(a^{\top}x)$ are a cornerstone of approximation theory, since they capture the simplest form of high-dimensional functions depending only on low-dimensional projections. They have been widely studied in numerical analysis, harmonic analysis, and machine learning (see, e.g., [pinkus2015ridge, barron1993universal]), both as hard test cases for approximation lower bounds and as building blocks for constructive approximation schemes. In practice, ridge-like structures appear in regression with latent variables, signal demixing, and physics-informed learning where dominant directions of variation exist.
@@ 2 Periodic CNN Architecture
Before analyzing approximation properties, we carefully describe the class of networks under study. The periodic CNN is obtained by replacing the usual finite-support convolutional operators with their circular analogues, i.e., convolutions performed under periodic boundary conditions. This modification equips the network with a natural translation-equivariant structure on the torus $\mathbb{T}^{d}$, aligning the architecture with many scientific settings where data are inherently periodic (e.g., spectral signals, crystallographic structures, or functions defined on wrapped domains).
@@ 2.1 Prelims
In this section, we introduce the deep CNNs structure with its mathematical definition.
@@ 3.1 Approximation of (d1)(d-1)-dimensional ridge functions
We proceed the proof in the following steps
@@ 3.2 Impossibility for (d2)(d-2)-dimensional ridge functions
Consequently the Fourier coefficients of $f$ are supported on the one-dimensional lattice
@@
Consequently the linear action of $J$ convolutional layers satisfies

### v2

@@ 1 Introduction
In this work, we focus on a new architectural variant we call the periodic CNN. The key modification is the enforcement of periodic boundary conditions in convolutional operations. Such a structure arises naturally in many scientific and engineering applications, for example, when signals live on tori, in materials with repeating microstructure, or in spectral data analysis.
@@
This dichotomy yields both theoretical insights and practical implications. On the theory side, it provides one of the first precise separation results for CNN variants, showing exactly where periodicity enables expressivity and where it constrains it. On the practical side, it suggests that periodic CNNs are particularly well matched to high-dimensional structured signals where information is concentrated along d1d-1 directions, a common scenario in physics-informed learning and scientific computing.
@@
Ridge functions $f(x)=\phi(a^{\top}x)$ are a cornerstone of approximation theory, since they capture the simplest form of high-dimensional functions depending only on low-dimensional projections. They have been widely studied in numerical analysis, harmonic analysis, and machine learning (see, e.g., [pinkus2015ridge, barron1993universal]), both as hard test cases for approximation lower bounds and as building blocks for constructive approximation schemes. In practice, ridge-shaped structures appear in regression with latent variables, signal remixing, and physics-informed learning where dominant directions of variation exist.
@@ 2 Periodic CNN Architecture
Before analyzing approximation properties, we carefully describe the class of networks under study. The periodic CNN is obtained by replacing the usual finite-support convolutional operators with their circular analogs, i.e., convolutions performed under periodic boundary conditions. This modification equips the network with a natural translation-equivariant structure on the torus $\mathbb{T}^{d}$, aligning the architecture with many scientific settings where data are inherently periodic (e.g., spectral signals, crystallographic structures, or functions defined on wrapped domains).
@@ 2.1 Prelims
In this section, we introduce the deep CNN structure with its mathematical definition.
@@ 3.1 Approximation of (d1)(d-1)-dimensional ridge functions
We proceed with the proof in the following steps
@@ 3.2 Impossibility for (d2)(d-2)-dimensional ridge functions
Consequently, the Fourier coefficients of $f$ are supported on the one-dimensional lattice
@@
Consequently, the linear action of $J$ convolutional layers satisfies